OpenStudy (anonymous):

A function is shown below: f(x) = x^3 + 5x^2 - x - 5 Part A: What are the factors of f(x)? Show your work. Part B: What are the zeros of f(x)? Show your work. Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

OpenStudy (anonymous):

@UnkleRhaukus may you please help me? I really need to get this done and id love for you to help!

OpenStudy (unklerhaukus):

\[f(x)=x^3 + 5x^2 - x - 5\] look at the first two terms \[x^3 \text{ and } 5x^2\] what is a common factors of these?

OpenStudy (unklerhaukus):

are you there ? @Floridagirl10

OpenStudy (anonymous):

yes sorry this website is slow for me... in order to find the common factor of those two.. wouldn't x^@ be 1x^2 ? the common factor would be 1...? @UnkleRhaukus

OpenStudy (unklerhaukus):

yeah openstudy is having some issues at the moment, i'm not quite sure what you'r saying, are you saying the common of x^3 and 5x^2 is x^2 ?

OpenStudy (anonymous):

I honestly don't know im lost right now.. Sorry. Well I think I am saying that the common factor is x^2. when im typing it doesn't come up til after I send it so I ciouldnt see it.

OpenStudy (unklerhaukus):

ok well in that case you are correct! the common factor is x^2 This means we can write \[x^3+5x^2\]as\[x^2(\,\cdots\,+\,\cdots\,)\] [fill in the blanks]

OpenStudy (anonymous):

would it be x^2(x+5) ? @UnkleRhaukus

OpenStudy (unklerhaukus):

Excellent! so we have factored the equation a little \[f(x)= x^3 + 5x^2 - x - 5\\ \qquad=x^2(x+5)-x-5\]

OpenStudy (unklerhaukus):

Now lets look at the last two terms -x and -5 what is the common factor this time?

OpenStudy (anonymous):

Im not sure about this one.. :/

OpenStudy (unklerhaukus):

they are both negative right?

OpenStudy (unklerhaukus):

you could write -x as (-1)(x) and -5 as (-1)(5)

OpenStudy (unklerhaukus):

so the common factor here is -1 This means we can write -x -1 as -( ⋯ + ⋯ ) what are the blanks?

OpenStudy (anonymous):

iit will be -(x+1)? or is it -1?

OpenStudy (unklerhaukus):

whoops i mean -x -5 becomes -( ⋯ + ⋯ )